OpenStudy (anonymous):

Solve the equation: 3/(x+1) - 1/x = 1/2

OpenStudy (anonymous):

omg i forgot how to do dis @blondie16 <-- pro

OpenStudy (anonymous):

\[\frac{ 3 }{ (x+1) }-\frac{ 1 }{ x }=\frac{ 1 }{ 2 }\] ^^that what is supposed to look like?^^

OpenStudy (anonymous):

so I put 1/2 as the answer?

OpenStudy (anonymous):

im asking if thats what the question is supposed to look like lol

OpenStudy (anonymous):

lol I have no idea, how I put the question is how it looks

OpenStudy (anonymous):

alright hmmmmm well you would have to solve the equation for x

OpenStudy (anonymous):

And hows that ? :o

hero (hero):